Mathematics 133 – Intensive Calculus for Science 1 Quiz 7 – October 28, 2005 Your Name: Do all work on this sheet (front and back if needed). You have 15 minutes. Calculators allowed on this quiz. Compute f′(x) for each of the following (you do not need to simplify) A) f(x) = 3 · 12^{x} − 2 · e^{x} + 2

B) f(x) = \frac{x−1}{x^{2}+4}

C) f(x) = (x^{4} + 6x^{2} + 9)^{23}

D) f(x) = e^{x+3}(x − \sqrt{x})